OpenStudy (anonymous):

If y = x + 5 were changed to y = x + 9, how would the graph of the new function compare with the first one? A. It would be steeper. B. It would be shifted right. C. It would be shifted up. D. It would be shifted down.

OpenStudy (michele_laino):

If I call \(f(x)=x+5\) and \(g(x)=x+9\), then I can write: \(g(x)=f(x)+4\)

OpenStudy (anonymous):

so ti would be c ?

OpenStudy (anonymous):

i'm confused

OpenStudy (michele_laino):

hint: 

OpenStudy (michele_laino):

correct! it is option C.

OpenStudy (anonymous):

thank you !!